OpenStudy (anonymous):

2x+5=50-4x Solve for x pleasee help

OpenStudy (phoenixfire):

Move all the terms with x to one side of the equation. And all terms without an x to the other side. What have you got?

OpenStudy (anonymous):

2x=3x=50+5???

OpenStudy (phoenixfire):

When you move a term from one side to the other side you have to change the sign. Here's the proper method: \[a+b=c\]To isolate b we subtract a from both sides of the equation. \[b+a-a=c-a\]You can see that a-a=0 leaving you with\[b=c-a\] A simpler way to remember is to change the sign if you move a term to the other side. \[2x+5=50-4x\]Move the -4x to the left. and the +5 to the right.\[2x+4x=50-5\] Can you solve from here?

OpenStudy (anonymous):

umm yeah i can thanks

OpenStudy (anonymous):

ummm actually oi tried and my final answer wasnt correct

OpenStudy (phoenixfire):

Well show me how you've done it. Each step.

OpenStudy (anonymous):

2x+4x=50-5 6x=45 - - 6 6 x=7.5 WRONG the final answer is supposed to be 20

OpenStudy (phoenixfire):

Your answer book is wrong. Or you are looking at the wrong answer. x=7.5 is correct if you are solving for x.

OpenStudy (anonymous):

should i attach the picture of this problem? so you can see?

OpenStudy (phoenixfire):

Yes please.

OpenStudy (anonymous):

number 7 @PhoenixFire

OpenStudy (anonymous):

ohhhhh i know i didnt substitute it in the problem lol my badddd

OpenStudy (phoenixfire):

So you got it?

OpenStudy (anonymous):

yeah lol thanks